Soo bought 6 CDs that were each the same price. Including sales tax, she paid a total of $71.40.
Each CD had a tax of $1.10.
What was the price of each CD before tax?

Answer by stanbon(75887)

Soo bought 6 CDs that were each the same price. Including sales tax, she paid a total of $71.40.
Each CD had a tax of $1.10.
What was the price of each CD before tax?
----
Price of 6 CD's before tax = 71.40 - 6*110 = $64.80
---
Price of each CD before tax = 64.80/6 = $10.80
--------------
Cheers,
Stan H.
----------